# Thinking in Space: How Multimodal Large Language Models See, Remember, and Recall Spaces

### v1

@@ 7 Related Works
Apart from visual-spatial intelligence in Sec.˜2, we further ground our work in the following two related areas:
MLLMs with Visual-Spatial Awareness. Building on the powerful language and reasoning abilities of LLMs [65, 66, 9, 79, 80, 3, 75] and the feature extraction abilities of modern vision encoders [29, 67, 63], MLLMs, especially visual MLLMs, exhibit unprecedented visual understanding capabilities [33, 76, 39, 83, 99, 88], a promising direction toward developing world models [48] and embodied agents [21, 57, 36, 17]. However, grounding MLLMs in the real world presents significant challenges for models’ visual-spatial intelligence, motivating recent efforts [91, 13, 16, 10, 46, 40, 102]. Unlike prior works, which primarily focus on understanding spatial information through 2D images [74, 68, 90] or solely language [87, 70, 89, 56, 87], our work assesses models’ visual spatial intelligence using real-world videos, which more closely mirrors human understanding of the world and application scenarios for embodied agents.
Benchmarking MLLMs on Video. With MLLMs displaying impressive performance on still-images across perception, reasoning, and multi-disciplinary tasks [96, 50, 38, 95], there is increasing interest in evaluating MLLMs’ video understanding capabilities [24, 53, 42, 61, 49, 43, 23, 93, 52]. For example, Video-MME [24] comprehensively evaluates MLLMs across various video-related tasks, including recognition and perception. EgoSchema [53] and OpenEQA [62] evaluate MLLMs’ understanding abilities using egocentric videos. Despite their significance, most prior works focus on content-level understanding [24, 42, 53, 61], which primarily serves as a temporal extension of 2D image understanding without 3D spatial consideration. Extending beyond prior benchmarks, our work establishes a testbed evaluating models’ 3D video-based visual-spatial intelligence, using video as an interface to understand the real world.
@@ 8 Discussion and Future Work
Acknowledgements.
We thank Ellis Brown, Ryan Inkook Chun, Youming Deng, Oscar Michel, Srivats Poddar, Xichen Pan, Austin Wang, Gavin Yang, and Boyang Zheng for their contributions as human annotators and evaluators. We also thank Fred Lu for proofreading our manuscript. We also thank Chen Feng, Richard Tucker, Noah Snavely, Leo Guibas, and Rob Fergus for their helpful discussions and feedback. This work was mainly supported by the Open Path AI Foundation, Google TPU Research Cloud (TRC) program, and the Google Cloud Research Credits program (GCP19980904). S.X. thanks the OpenAI Researcher Access program and an Amazon Research award for their support.
@@ Appendix A Appendix Outline
  • Technical details about VSI-Bench construction and our linguistic and visual analysis (Appendix˜B);
  • Evaluation setup and full evaluation results for VSI-Bench sub-experiments (Appendix˜C);
  • Analysis on input sequencing and repetition (Appendix˜D);
  • Additional visualization results (Appendix˜E).
@@ B.1 VSI-Bench Construction Pipeline
The question templates for the generation of each task are listed in Tab.˜3.
@@ B.3 Implementation Details of CoT Methods
Fig.˜16. Fig.˜17, and Fig.˜18 illustrate these three prompting techniques and model outputs under the different strategies.
@@ B.4 Cognitive Map
For the categories of interest, we include all potential categories as shown in Fig.˜11 and Fig.˜12. Such setup facilitates our focus on assessing the spatial awareness of the MLLM rather than its perceptual capabilities. In contrast, for benchmark tasks such as evaluating relative distance (as shown in Tab.˜2), we restrict the provided categories to those explicitly mentioned in each question. This ensures that no additional information apart from the question is included.
Distance Locality Calculation. To quantitatively evaluate the cognitive maps, we measure inter-category distances as illustrated in Fig.˜12. Specifically, for each category, we compute its Euclidean distance to all other categories. When a category contains multiple objects, we define the inter-category distance as the shortest distance between any two objects from the respective categories. We perform these distance calculations on both MLLM-predicted and ground truth cognitive maps and consider an MLLM’s predicted distance between two categories to be correct if it differs from the ground truth distance by no more than one grid unit. We apply this evaluation process across all cognitive maps and group the distances into eight bins to calculate the average accuracy on different bins.
@@ C.1 General Evaluation Setup
Our evaluation processes are primarily conducted using the LMMs-Eval project [97]. To ensure reproducibility, unless otherwise specified, we adopt a greedy decoding strategy for all models (i.e., the temperature is set to 0, and both top-p and top-k are set to 1). The input for the models is formatted as follows: [Video Frames][Pre-prompt][Question][Post-prompt], where Question includes the question and any available options. The specific Pre-prompt and Post-prompt for different models and question types are detailed in Tab.˜6.
@@ C.3 Number of Frames Setup
Typically, MLLMs subsample a fixed number of frames for evaluation. For all open-source models and the GPT-4 API, following [97], we manually sample video frames from the entire video at evenly spaced time intervals. For the Gemini API, we follow its instructions, uploading and feeding the entire video to the model. The number of frames used for each model are provided in Tab.˜5.
@@ C.4 More Evaluation Results
Here, we provide more evaluation results on our benchmark, including the full evaluation results of VSI-Bench (tiny), blind evaluation results, and vision-enabled - vision-disabled results.
VSI-Bench (tiny) Results. As shown in Tab.˜7, we provide the evaluation results of all models on VSI-Bench (tiny). The rankings and average accuracy of MLLMs on VSI-Bench (tiny) remain consistent to the results reported in Fig.˜6. This consistency suggests that the human evaluation and analysis results conducted on VSI-Bench (tiny) are reliable.
Blind Evaluation. As shown in Tab.˜8, we present the evaluation results for all MLLMs on VSI-Bench. Generally, larger variants within the same model family often demonstrate better performance in blind evaluations, as seen in comparisons such as Gemini-1.5 Flash vs. Gemini-1.5 Pro and VILA-1.5-8B vs. VILA-1.5-40B. The blind evaluation also highlights LLM biases across tasks. For instance, LongVILA-8B achieves 47.5% accuracy on the object count task, benefiting from a bias that frequently leads it to predict 2 as the answer.
Vision Enabled - Vision Disabled. Tab.˜9 presents the improvement of MLLMs from using visual signals to answer VSI-Bench. Almost all MLLMs obtain improvements from visual signals, with notable improvements in tasks such as object count, room size, relative distance and appearance order.
@@ Appendix D Input Sequencing and Repetition Analysis
MLLM’s performance degrades with question-first paradigm. As shown in Tab.˜4 (a), switching to a video-first approach results in a 2.5% decrease in overall performance for Gemini compared to the question-first approach.
@@
MLLM benefits from multiple video views. In addition, humans often improve their VQA performance by reviewing visual content multiple times, inspiring us to implement a similar setup for MLLMs. As shown in Tab.˜4 (b), Gemini achieves a notable 2.1% performance gain with two repeated videos as input. This is surprising, as autoregressive MLLMs theoretically have the capability to revisit the video multiple times during answer generation, even if the video is only presented once. This finding suggests that, despite its remarkable capabilities, a powerful MLLM like Gemini still has suboptimal reasoning processes for Video QA.
@@ E.1 VSI-Bench Examples
In Fig.˜13 and Fig.˜14, we provide more examples from VSI-Bench to illustrate the structure and format of tasks, questions, and answers.
@@ E.2 Error Analysis Examples
In Fig.˜15, we present more case studies for our human-conducted error analysis on VSI-Bench. In the error analysis, we identify the categorized error types and highlight the relevant parts of the explanation.
@@
InternVL2-2B 8
InternVL2-8B 8
InternVL2-40B 8
@@
LLaVA-NeXT-Video-7B 32
LLaVA-NeXT-Video-72B 32
@@ E.3 Linguistic Prompting Examples
We provide examples for the three CoT prompting methods discussed in Sec.˜5.2 to illustrate their concrete reasoning procedure in detail. We include examples of three selected tasks: object count, object size, and room size. For Zero-Shot Chain of Thought, as shown in Fig.˜16, we highlight each step of the MLLM’s reasoning process to offer insights into how it arrives at its final decision. For Self-Consistency w/ CoT, as illustrated in Fig.˜17, each example is paired with five independent responses. The final answer is then determined by a majority vote. For Tree-of-Thought, Fig.˜18 details how each depth of the decision tree is reached. At the first depth, the MLLM generates three potential plans and conducts a choice analysis to select the optimal plan. At the second and final depth, the selected plan is used to generate three potential answers, with the final output determined through a majority vote.
@@ E.4 Cognitive Map Examples
In Fig.˜19, we include 10 additional cognitive maps and pair each prediction with its corresponding ground truth map to provide insight into the alignment between predicted and ground truth layouts.
@@
LLaVA-NeXT-Video-7B 35.7 49.0 12.8 48.6 21.4 40.0 43.5 34.0 36.0
LLaVA-NeXT-Video-72B 39.3 41.4 26.6 55.6 31.6 36.0 25.6 42.0 56.0
@@
LLaVA-NeXT-Video-7B 25.2 14.8 14.6 32.5 26.1 26.8 45.0 33.0 8.5
LLaVA-NeXT-Video-72B 29.1 19.0 25.4 46.3 26.1 29.0 38.8 33.0 15.5
@@
InternVL2-2B 9.6 16.4 1.2 12.8 35.0 6.9 3.0 2.5 -0.8
InternVL2-8B 7.0 -8.8 1.9 9.9 39.1 9.7 -8.5 -3.0 16.0
InternVL2-40B 11.6 29.6 -2.2 7.3 31.1 11.8 -5.5 6.1 14.9
@@
LLaVA-NeXT-Video-7B 10.5 33.8 -0.6 15.2 -1.9 16.7 -2.7 1.0 22.1
LLaVA-NeXT-Video-72B 11.7 29.9 -2.6 11.1 9.2 13.3 -2.0 2.0 33.0
@@
Table 9: Results of Vision Enabled - Vision Disabled.

### v2

@@ 7 Related Works
Apart from visual-spatial intelligence in Sec. 2, we further ground our work in the following two related areas:
MLLMs with Visual-Spatial Awareness. Building on the powerful language and reasoning abilities of LLMs [67, 68, 9, 81, 82, 3, 77] and the feature extraction abilities of modern vision encoders [30, 69, 65], MLLMs, especially visual MLLMs, exhibit unprecedented visual understanding capabilities [34, 78, 40, 86, 103, 91, 102], a promising direction toward developing world models [50] and embodied agents [21, 59, 37, 17]. However, grounding MLLMs in the real world presents significant challenges for models’ visual-spatial intelligence, motivating recent efforts [94, 13, 16, 10, 48, 41, 107, 29]. Unlike prior works, which primarily focus on understanding spatial information through 2D images [76, 70, 93] or solely language [90, 72, 92, 58, 90], our work assesses models’ visual spatial intelligence using real-world videos, which more closely mirrors human understanding of the world and application scenarios for embodied agents.
Benchmarking MLLMs on Video. With MLLMs displaying impressive performance on still-images across perception, reasoning, and multi-disciplinary tasks [99, 52, 39, 98], there is increasing interest in evaluating MLLMs’ video understanding capabilities [24, 55, 43, 63, 51, 44, 23, 96, 54, 47, 85]. For example, Video-MME [24] comprehensively evaluates MLLMs across various video-related tasks, including recognition and perception. EgoSchema [55] and OpenEQA [64] evaluate MLLMs’ understanding abilities using egocentric videos. Despite their significance, most prior works focus on content-level understanding [24, 43, 55, 63], which primarily serves as a temporal extension of 2D image understanding without 3D spatial consideration. Extending beyond prior benchmarks, our work focuses on spatial intelligence and requires core spatial capabilities like visual working memory and implicit scene reconstruction.
@@ 8 Discussion and Future Work
Acknowledgments. We thank Ellis Brown, Ryan Inkook Chun, Youming Deng, Oscar Michel, Srivats Poddar, Xichen Pan, Austin Wang, Gavin Yang, and Boyang Zheng for their contributions as human annotators and evaluators. We also thank Fred Lu for proofreading our manuscript. We also thank Chen Feng, Richard Tucker, Noah Snavely, Leo Guibas, and Rob Fergus for their helpful discussions and feedback. This work was mainly supported by the Open Path AI Foundation, Google TPU Research Cloud (TRC) program, and the Google Cloud Research Credits program (GCP19980904). SX also acknowledges support from Intel AI SRS, IITP grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project), Amazon Research Award, and NSF Award IIS-2443404.
@@ Appendix A Appendix Outline
  • Technical details about VSI-Bench construction and our linguistic and visual analysis (Appendix B);
  • Evaluation setup and full evaluation results for VSI-Bench sub-experiments (Appendix C);
  • Analysis on input sequencing and repetition (Appendix D);
  • Additional visualization results (Appendix E).
@@ B.1 VSI-Bench Construction Pipeline
The question templates for the generation of each task are listed in Tab. 3.
@@ B.3 Implementation Details of CoT Methods
Fig. 17. Fig. 18, and Fig. 19 illustrate these three prompting techniques and model outputs under the different strategies.
@@ B.4 Cognitive Map
For the categories of interest, we include all potential categories as shown in Fig. 10 and Fig. 11. Such setup facilitates our focus on assessing the spatial awareness of the MLLM rather than its perceptual capabilities. In contrast, for benchmark tasks such as evaluating relative distance (as shown in Tab. 2), we restrict the provided categories to those explicitly mentioned in each question. This ensures that no additional information apart from the question is included.
Distance Locality Calculation. To quantitatively evaluate the cognitive maps, we measure inter-category distances as illustrated in Fig. 11. Specifically, for each category, we compute its Euclidean distance to all other categories. When a category contains multiple objects, we define the inter-category distance as the shortest distance between any two objects from the respective categories. We perform these distance calculations on both MLLM-predicted and ground truth cognitive maps and consider an MLLM’s predicted distance between two categories to be correct if it differs from the ground truth distance by no more than one grid unit. We apply this evaluation process across all cognitive maps and group the distances into eight bins to calculate the average accuracy on different bins.
B.5 Cognitive Map on More MLLMs
We evaluate two more MLLMs, LLaVA-Video-7B and LLaVA-Video-72B. Tab. 4 validates our Sec. 6.1 finding of significantly stronger local than global accuracy. Regarding Sec. 6.2, as shown in Tab. 5, LLaVA-Video-72B achieves an 8% performance gain. In contrast, LLaVA-Video-7B performance decreases, likely due to its limited model capacity, which impairs cog. map prediction (Tab. 4 shows its suboptimal acc. on cog. map compared to Gemini-1.5 Pro and LLaVA-Video-72B).
Distance [1.0, 2.1] (2.1, 3.3] (3.3, 4.4] (4.4, 5.5] (5.5, 6.6] (6.6, 7.8] (7.8, 8.9] (8.9, 10.0]
Gemini-1.5 Pro 0.64 0.48 0.35 0.35 0.28 0.12 0.06 0.00
LLaVA-Video-72B 0.59 0.45 0.42 0.30 0.15 0.23 0.16 0.00
LLaVA-Video-7B 0.50 0.43 0.34 0.29 0.19 0.18 0.14 0.00
Table 4: Locality of cognitive maps.
Models LLaVA-Video-72B LLaVA-Video-7B
w/o. Cog. Map 36.0 40.0
w/.   Cog. Map 42.0 32.0
Table 5: Rel. dist. task with cognitive maps.
@@ C.1 General Evaluation Setup
Our evaluation processes are primarily conducted using the LMMs-Eval project [100]. To ensure reproducibility, unless otherwise specified, we adopt a greedy decoding strategy for all models (i.e., the temperature is set to 0, and both top-p and top-k are set to 1). The input for the models is formatted as follows: [Video Frames][Pre-prompt][Question][Post-prompt], where Question includes the question and any available options. The specific Pre-prompt and Post-prompt for different models and question types are detailed in Tab. 9.
@@ C.2 Human Evaluation Setup
In addition, we provide the human evaluation on another VSI-Bench subset with 560 samples optimized to minimize the average performance gap between this subset and full set for all MLLMs. As shown in Tab. 10, this subset has an average performance discrepancy compared to full set (see Fig. 6) just 0.5% and a maximum of 2.9%.
@@ C.3 Number of Frames Setup
Typically, MLLMs subsample a fixed number of frames for evaluation. For all open-source models and the GPT-4 API, following [100], we manually sample video frames from the entire video at evenly spaced time intervals. For the Gemini API, we follow its instructions, uploading and feeding the entire video to the model. The number of frames used for each model are provided in Tab. 8. We believe that frame sampling strategies are a model design choice separate from the benchmark design. Established benchmarks (e.g., VideoMME [24] and EgoSchema [55]) also employ default sampling, reinforcing this perspective. In addition, as shown in the Fig. 12, the # of sampled frames only marginally affects performance—it is not the primary bottleneck.
Figure 12: Analysis of different # sampled frames.
@@ C.4 More Evaluation Results
Here, we provide more evaluation results on our benchmark, including blind evaluation results, the Socratic LLMs, the full evaluation results of VSI-Bench (tiny), and vision-enabled -- vision-disabled results.
Figure 13: Performance comparisons between Vision Enabled (w/ video), Vision Disabled (w/o video) and Chance Level (Freq.). Enabled--Disabled indicates the gap between Vision Enabled and Vision Disabled, and Disabled--Chance betokens the gap between Vision Disabled and Chance Level (Freq.). Tasks are sorted by Enable--Disable for better understanding.
Blind Evaluation. We compare MLLMs’ performance against “Chance Level (frequency)” and “Vision Disabled” (blind) results, using averages across six of the strongest models (3 open-source and 3 closed-source). As shown in Fig. 13, the consistent improvements in “Enabled--Disabled” and general degradation in “Disabled--Chance” demonstrates that video is essential and beneficial for our VSI-Bench, with blind models performing below chance level. Meanwhile, MLLMs struggle to improve beyond chance level in the absolute distance estimation, route plan, and relative direction tasks, whether vision is enabled or not, underscoring the difficulty of these tasks. Note that on object size, “Vision Disabled” models already significantly outperform chance level, likely due to common-sense knowledge learned during language model training.
In addition, as shown in Tab. 12, we present the evaluation results for all MLLMs on VSI-Bench. Generally, larger variants within the same model family often demonstrate better performance in blind evaluations, as seen in comparisons such as Gemini-1.5 Flash vs. Gemini-1.5 Pro and VILA-1.5-8B vs. VILA-1.5-40B. The blind evaluation also highlights LLM biases across tasks. For instance, LongVILA-8B achieves 47.5% accuracy on the object count task, benefiting from a bias that frequently leads it to predict 2 as the answer.
Socratic LLMs with Frame Captions. Following OpenEQA [54] and HourVideo [12], we implement a Socratic variant of GPT-4o using LLaVA-Video-72B as the captioner and GPT-4o as the answering LLM. As shown in Tab. 6, Socratic lags behind the standard GPT-4o by 4.7%.
GPT-4o Standard Socratic Blind
Avg. 34.0 29.3 14.5
Table 6: Socratic LLMs with Frame Captions.
VSI-Bench (tiny) Results. As shown in Tab. 11, we provide the evaluation results of all models on VSI-Bench (tiny). The rankings and average accuracy of MLLMs on VSI-Bench (tiny) remain consistent to the results reported in Fig. 6. This consistency suggests that the human evaluation and analysis results conducted on VSI-Bench (tiny) are reliable.
Vision Enabled -- Vision Disabled. Tab. 13 presents the improvement of MLLMs from using visual signals to answer VSI-Bench. Almost all MLLMs obtain improvements from visual signals, with notable improvements in tasks such as object count, room size, relative distance and appearance order.
@@ Appendix D Input Sequencing and Repetition Analysis
MLLM’s performance degrades with question-first paradigm. As shown in Tab. 7 (a), switching to a video-first approach results in a 2.5% decrease in overall performance for Gemini compared to the question-first approach.
@@
MLLM benefits from multiple video views. In addition, humans often improve their VQA performance by reviewing visual content multiple times, inspiring us to implement a similar setup for MLLMs. Specifically, input is formatted as: [Video] [Context] [Video] with identical video, where the system prompt explicitly informs the model of the redundancy of input video. As shown in Tab. 7 (b), Gemini achieves a notable 2.1% performance gain with two repeated videos as input. This is surprising, as autoregressive MLLMs theoretically have the capability to revisit the video multiple times during answer generation, even if the video is only presented once. This finding suggests that, despite its remarkable capabilities, a powerful MLLM like Gemini still has suboptimal reasoning processes for Video QA.
@@ E.1 VSI-Bench Examples
In Fig. 14 and Fig. 15, we provide more examples from VSI-Bench to illustrate the structure and format of tasks, questions, and answers.
@@ E.2 Error Analysis Examples
In Fig. 16, we present more case studies for our human-conducted error analysis on VSI-Bench. In the error analysis, we identify the categorized error types and highlight the relevant parts of the explanation.
@@
InternVL2-2B 32
InternVL2-8B 32
InternVL2-40B 32
@@
LLaVA-Video-7B 32
LLaVA-Video-72B 32
@@ E.3 Linguistic Prompting Examples
We provide examples for the three CoT prompting methods discussed in Sec. 5.2 to illustrate their concrete reasoning procedure in detail. We include examples of three selected tasks: object count, object size, and room size. For Zero-Shot Chain of Thought, as shown in Fig. 17, we highlight each step of the MLLM’s reasoning process to offer insights into how it arrives at its final decision. For Self-Consistency w/ CoT, as illustrated in Fig. 18, each example is paired with five independent responses. The final answer is then determined by a majority vote. For Tree-of-Thought, Fig. 19 details how each depth of the decision tree is reached. At the first depth, the MLLM generates three potential plans and conducts a choice analysis to select the optimal plan. At the second and final depth, the selected plan is used to generate three potential answers, with the final output determined through a majority vote.
@@ E.4 Cognitive Map Examples
In Fig. 20, we include 10 additional cognitive maps and pair each prediction with its corresponding ground truth map to provide insight into the alignment between predicted and ground truth layouts.
@@
Obj. Count
Abs. Dist.
Obj. Size
Room Size
Rel. Dist.
Rel. Dir.
Route Plan
Appr. Order
Methods Avg. Numerical Answer Multiple-Choice Answer
Proprietary Models (API)
Gemini-1.5 Flash 41.6 49.1 30.3 52.7 53.7 37.1 40.8 31.4 37.1
Gemini-1.5 Pro 44.9 55.1 30.3 63.1 43.3 50.0 45.9 35.7 35.7
Open-source Models
InternVL2-2B 27.0 22.4 24.9 21.1 34.1 32.9 43.5 30.0 7.1
InternVL2-8B 34.1 22.6 28.3 47.6 39.6 35.7 30.4 30.0 38.6
InternVL2-40B 35.5 34.4 26.9 45.6 31.3 41.4 31.7 32.9 40.0
LongVILA-8B 21.0 28.7 8.6 16.3 0.0 28.6 30.5 31.4 24.3
VILA-1.5-8B 28.4 17.3 21.6 49.9 18.6 31.4 34.4 30.0 24.3
VILA-1.5-40B 30.8 21.4 24.4 48.3 21.9 40.0 25.0 30.0 35.7
LongVA-7B 29.0 38.1 16.9 38.1 21.7 32.9 42.8 25.7 15.7
LLaVA-Video-7B 34.9 47.9 13.4 46.7 23.9 42.9 41.9 32.9 30.0
LLaVA-Video-72B 40.5 48.3 22.6 56.7 34.6 41.4 36.5 35.7 48.6
LLaVA-OneVision-0.5B 27.6 45.1 27.9 14.7 27.9 28.6 37.0 34.3 5.7
LLaVA-OneVision-7B 32.1 46.9 19.9 46.9 12.1 41.4 35.1 30.0 24.3
LLaVA-OneVision-72B 39.6 42.7 23.7 56.7 36.9 41.4 39.5 31.4 44.3
Table 10: Evaluation results on VSI-Bench 560 samples subset.
@@
LLaVA-Video-7B 35.7 49.0 12.8 48.6 21.4 40.0 43.5 34.0 36.0
LLaVA-Video-72B 39.3 41.4 26.6 55.6 31.6 36.0 25.6 42.0 56.0
@@
LLaVA-Video-7B 25.2 14.8 14.6 32.5 26.1 26.8 45.0 33.0 8.5
LLaVA-Video-72B 29.1 19.0 25.4 46.3 26.1 29.0 38.8 33.0 15.5
@@
InternVL2-2B 8.7 20.3 0.3 10.8 29.2 5.2 2.9 2.5 -1.6
InternVL2-8B 9.9 -0.6 2.2 10.6 43.5 10.9 -5.8 -4.1 22.8
InternVL2-40B 12.6 35.9 -2.9 9.0 26.8 17.3 -5.0 9.9 20.0
@@
LLaVA-Video-7B 10.5 33.8 -0.6 15.2 -1.9 16.7 -2.7 1.0 22.1
LLaVA-Video-72B 11.7 29.9 -2.6 11.1 9.2 13.3 -2.0 2.0 33.0
@@
Table 13: Results of Vision Enabled -- Vision Disabled.